OpenStudy (anonymous):

4/5(3x+4) less than equal to 20. Please help! I don't know how to get rid of the fraction before PEMDAS applies.

OpenStudy (anonymous):

I suggest you use distributive property.

OpenStudy (anonymous):

Distribute the 5 into the 3x and 4.

OpenStudy (anonymous):

Is the inequality this: \[\frac{4}{5}(3x+4) \le 20\]or\[\frac{4}{5(3x+4) }\le 20\]

OpenStudy (anonymous):

4/5(3x+4)≤20

OpenStudy (anonymous):

repeating the same thing doesn't clarify anything.

OpenStudy (anonymous):

Sorry this is my first time using this. If I distribute the 5 into 3x and 4 and 20 I get \[4(15x+20)\le100\]

OpenStudy (anonymous):

Then...distribute 4 ... \[60x+80\le20 \]

OpenStudy (anonymous):

The answer in the text book says it is -7 and I have no idea how that answer can apply.

OpenStudy (anonymous):

Ok if the answer is \(x \le 7\) then it must be: \[\frac{4}{5}(3x + 4) \le 20\]

OpenStudy (anonymous):

You multiply both sides by 5. And you have: \[4(3x+4) \le 100\]

OpenStudy (anonymous):

Then continue from there.